I need help with rearranging an equation. Can you please show a step by step method of how to get to the S value. Thank you.
[S/(45,000,000 + S)] x 20,412,000= 882,000

Found 2 solutions by Theo, Alan3354:
Answer by Theo(13342)

(S/(45,000,000 + S) * 20,412,000 = 882,000

multiply both sides of the equation by (45,000,000 + S) to get:

S * 20,412,000 = 882,000 * (45,000,000 + S)

simplify by distributing the multiplication on the right side of the equation to get:

S * 20,412,000 = 882,000 * 45,000,000 + 882,000 * S

subtract 882,000 * S from both sides of the equation to get:

S * 20,412,000 - 882,000 * S = 882,000 * 45,000,000

factor out the S on the left side of the equation to get S * (20,412,000 - 882,000) = 882,000 * 45,000,000

divide both sides of the equation by (20,412,000 - 882,000) to get S = (882,000 * 45,000,000) / (20,412,000 - 882,000)

solve for S to get S = 2032258.065

go back to your original equation and replace S with that and you will see that the original equation is true, confirming the solution is correct.





Answer by Alan3354(69443)

I need help with rearranging an equation. Can you please show a step by step method of how to get to the S value. Thank you.
[S/(45,000,000 + S)] * 20,412,000= 882,000
--- Use * (Shift 8) for multiply, not x.
-----
Divide by 20412000
[S/(45,000,000 + S)] = 882,000/20412000 = 882/20412 = 7/162
------
Multiply by 45000000+S
S = (7/162)*(45000000+S) = (3150000000 + 7S)/162
Multiply by 162
162S = 3150000000 + 7S
Subtract 7S
155S = 3150000000
Divide by 155